OpenStudy (anonymous):

(1)/(2)+(1)/(3)=(1)/(x)?

OpenStudy (anonymous):

\[\frac{1}{2}+\frac{1}{3}=\frac{1}{x}\] Now we want to do Low Common Denominator (LCD) on the left side. How can we do that?